OpenStudy (calculusxy):

Ashley saved a distance equal to 80% of the length of the shortest side of a rectangular field by cutting across the diagonal of the field instead of along two of the sides. Find the ratio of the length of the shortest side of the field to the length of its longest side.

OpenStudy (calculusxy):

@eliesaab

OpenStudy (calculusxy):

@Loser66

OpenStudy (mathmate):

As in all geometry problems, making a diagram is half the problem solved.  We know that the diagonal is shorter than (x+kx) by 80% of x. Can you put this in an equation?

OpenStudy (calculusxy):

This is the equation that I got \[(\sqrt{a^2+b^2})^2 = (b + 0.2a)^2\]

OpenStudy (mathmate):

That looks good! So a is the shorter side, right? Can you continue?

OpenStudy (calculusxy):

So I then got 0.96a^2 = 0.4ab after some steps

OpenStudy (mathmate):

So far so good! :) So what would you do next?

OpenStudy (calculusxy):

I thought about dividing both sides by 0.4a 0.24a = b

OpenStudy (calculusxy):

I meant 2.4a = b

OpenStudy (mathmate):

You're almost there, keep going @calculusxy

OpenStudy (calculusxy):

I assumed that the shorter side be a. So do I divide both sides by a? 2.4 = a/b?

OpenStudy (mathmate):

Ouch! `Find the ratio of the length of the shortest side of the field to the length of its longest side.`

OpenStudy (calculusxy):

That's where I am having trouble--understanding what variable to isolate...

OpenStudy (calculusxy):